OpenStudy (quinton1012):

Solve for x. -3x + 4 = -8 A) x = 4 B) x = -4 C) No Solution D) x = - 4 3

OpenStudy (chris__lopez):

i think it might be b

OpenStudy (quinton1012):

thx can you help me with one more?

OpenStudy (quinton1012):

Which equation has solution x = -3? A) 2x - 7 = -1 B) 3x + 8 = 1 C) 1 2 x + 8 = 10 D) 1 2 (2x - 6) = -6

OpenStudy (chris__lopez):

hang on a bit

OpenStudy (quinton1012):

Okay

OpenStudy (quinton1012):

btw the 1 is 1/2 2

OpenStudy (quinton1012):

@knov @YumYum247 @RhondaSommer

OpenStudy (quinton1012):

@XxStudyingStuffxX

OpenStudy (yumyum247):